IMPROVED ANALYSES OF INFRARED AND VISIBLE EMISSION SPECTRA OF THE H_{3} AND D_{3} MOLECULES

MICHEL VERVLOET, Laboratoire de Photophysique Moléculaire du CNRS, Bâtiment 350, Université de Paris-Sud, 91405 Orsay Cedex, France; \underline{JAMES K. G. WATSON}, Steacie Institute for Molecular Sciences, National Research Council of Canada, Ottawa, Ontario, Canada, K1A OR6.

Electronic emission spectra of the neutral H_{3} and D_{3} molecules obtained with a Droege-Engelking^{a} corona discharge source were described at a previous symposium.^{b} The present work is concerned with analyses of these data for the complex of 3d, 3s, and 3p_{0} upper states, where the symbols mean nl_{λ}. The lower states are 3p_{1} (for the infrared) and 2p_{0} (for the visible). In H_{3} a (3s, 3d_{0}) perturbation is recognised, in addition to the (3p_{0}, 3d_{1}) perturbation discussed previously.^{c} The decreased interference by H_{2} or D_{2} in these spectra has made it possible to improve the lower-state combination differences and constants. However, the standard deviations in the fits of the upper states of both H_{3} and D_{3} are still quite large. This is possibly due to numerous small perturbations by background states. It would also be desirable to replace the present effective Hamiltonians by a variational-type treatment of centrifugal effects.

aA. T. Droege and P. C. Engelking, Chem. Phys. Lett. 96, 316 (1983).

bPaper TA12 (1991).

cG. Herzberg, J. T. Hougen, and J. K. G. Watson, Can. J. Phys. 60, 1261 (1982)